# On the use of the Gram matrix for multivariate functional principal components analysis

### v1

@@ 1 Introduction
The key motivation of this paper is to investigate the duality between rows and columns of a data matrix to estimate the eigencomponents of a multivariate functional dataset. The duality between rows and columns of a data matrix is a fundamental concept in classical multivariate statistics (Escofier,, 1979; Saporta,, 1990). A data matrix typically represents a set of observations of multiple features, each row corresponds to an individual observation and each column corresponds to an individual feature. The duality between rows and columns refers to the fact that many statistical methodologies can be conducted either on the rows or the columns of the data matrix, and the results will be related to each other. For example, the principal components obtained from a PCA run on the rows of the data matrix are the same as the ones obtained from a PCA run on the columns of the matrix. The choice of method to use, based on criteria such as computational time or data storage needs, is thus left to the statistician. This concept has been widely studied in multivariate statistics (see, e.g., Pagès, (2014); Härdle and Simar, (2019)). In the context of functional data, this principle has received limited attention despite being mentioned in the seminal paper of FDA (Ramsay,, 1982). Ramsay and Silverman, (2005) briefly commented on it in a concluding remark of Chapter 8, while Kneip and Utikal, (2001) and Benko et al., (2009) utilized it to compute principal components for dense univariate functional data. Chen et al., (2017) also employ it to gain computational advantage when univariate functional data are sampled on a very dense grid. To the best of our knowledge, however, there is no available literature on its application to multivariate functional data that are observed on different dimensional domains. Our aim is therefore to investigate this duality for multivariate functional data observed on different dimensional domains and provide guidelines to statisticians on which method to use in different cases.
The remainder of the paper is organized as follows. In Section 2, we define multivariate functional data, with components that are observed on possibly different domains. In Section 3, we develop the duality between the observations’ space and the functional components’ space. The relationship between the eigencomponents of the covariance operator of the multivariate functional datasets and the eigencomponents of the inner-product matrix between the observations is derived in Section 4. Extensive simulations are given in Section 5. We also provide guidelines on which method to use with respect to different data characteristics. The paper concludes with a discussion and an outlook in Section 6.
@@ 2 Model
Let us consider a set of $N$ curves $\mathcal{X}=\{X_{1},\ldots,X_{n},\ldots,X_{N}\}$ generated as a random sample of the $P$-dimensional stochastic process $X$ with continuous trajectories. Unless otherwise stated, the data are assumed to be observed without error. The data can be viewed as a table with $N$ rows and $P$ columns where each entry is a curve, potentially on a multidimensional domain (see Figure 1). Each row of this matrix represents an observation; while each column represents a functional feature. At the intersection of row $n$ and column $p$, we thus have $X_{n}^{(p)}$ which is the curve that concerns the (functional) feature $p$ for the individual $n$.
@@
For $n\in\{1,\dots,N\}$, each observation $n$ is attributed the weight $\pi_{n}$ such that $\sum_{n}\pi_{n}=1$, e.g., $\pi_{n}=1/N$. For a given $p\in\{1,\dots,P\}$, the mean curve of the $p$th feature along the $N$ observations is denoted by $\mu^{(p)}$. This quantity can be computed as
$\mu^{(p)}(t_{p})=\sum_{n=1}^{N}\pi_{n}X_{n}^{(p)}(t_{p}),\quad t_{p}\in\mathcal{T}_{p},\quad p\in\{1,\dots,P\}.$
The cross-covariance function of the $p$th and $q$th features along the $N$ observations can be computed as, for $p=1,\dots,P$ and $q=1,\dots,P$,
$C_{p,q}(s_{p},t_{q})=\sum_{n=1}^{N}\pi_{n}X_{n}^{(p)}(s_{p})X_{n}^{(q)}(t_{q})-\mu^{(p)}(s_{p})\mu^{(q)}(t_{q}),\quad s_{p}\in\mathcal{T}_{p},\quad t_{q}\in\mathcal{T}_{q}.$ (5)
@@
For the set $\mathcal{X}$, the inner-product matrix, also called the Gram matrix, $\mathbf{M}$ is defined as a matrix of size $N\times N$ with entries
This matrix is symmetric, positive definite, and interpretable as a proximity matrix, each entry being the distance between the weighted observations.
2.1 Basis decomposition
In many practical situations, functional data are noisy and only observed at specific time points. To extract the underlying functional features of the data, smoothing and interpolation techniques are commonly employed. These techniques involve approximating the true underlying function generating the data by a finite-dimensional set of basis functions. Assume that for each feature $p=1,\dots,P$, there exists a set of basis of functions $\Psi^{(p)}=\{\psi_{k}^{(p)}\}_{1\leq k\leq K_{p}}$ such that each feature of each curve $n=1,\dots,N$ can be expanded using the basis:
$X_{n}^{(p)}(t_{p})=\sum_{k=1}^{K_{p}}c^{(p)}_{nk}\psi_{k}^{(p)}(t_{p}),\quad t_{p}\in\mathcal{T}_{p},$ (7)
where $\{c^{(p)}_{nk}\}_{1\leq k\leq K_{p}}$ is a set of coefficients for feature $p$ of observation $n$. We denote by $\overline{c}_{k}^{(p)}=\sum_{n=1}^{N}\pi_{n}c^{(p)}_{nk}$ the mean coefficient of feature $p$ corresponding to the $k$th basis function. The $p$th feature of the mean function can be then expanded in the same basis as:
$\mu^{(p)}(t_{p})=\sum_{k=1}^{K_{p}}\overline{c}_{k}^{(p)}\psi_{k}^{(p)}(t_{p}),\quad t_{p}\in\mathcal{T}_{p}.$ (8)
Similarly, the covariance function of the $p$th and $q$th features is given by:
$C_{p,q}(s_{p},t_{q})=\sum_{k=1}^{K_{p}}\sum_{l=1}^{K_{q}}\left(\sum_{n=1}^{N}\pi_{n}c^{(p)}_{nk}c^{(q)}_{nl}-\overline{c}_{k}^{(p)}\overline{c}_{l}^{(q)}\right)\psi_{k}^{(p)}(s_{p})\psi_{l}^{(q)}(t_{q}),\quad s_{p}\in\mathcal{T}_{p},\quad t_{q}\in\mathcal{T}_{q}.$ (9)
[…]
$\mathbf{C}=\begin{pmatrix}\mathbf{C}^{(1)}&\cdots&\mathbf{C}^{(P)}\\ \end{pmatrix},\quad\text{and}\quad\Psi(\mathbf{t})=\text{diag}\{\Psi^{(1)}(t_{1}),\dots,\Psi^{(P)}(t_{P})\},$ (10)
where
$\mathbf{C}^{(p)}=\begin{pmatrix}c^{(p)}_{11}&\cdots&c^{(p)}_{1K_{p}}\\ \vdots&\ddots&\vdots\\ c^{(p)}_{N1}&\cdots&c^{(p)}_{NK_{p}}\end{pmatrix}\\ \quad\text{and}\quad\Psi^{(p)}(t_{p})=\begin{pmatrix}\psi_{1}^{(p)}(t_{p})\\ \vdots\\ \psi_{K_{p}}^{(p)}(t_{p})\end{pmatrix}.$ (11)
Using the basis expansion and denoting $\Pi^{\top}=(\pi_{1},\dots,\pi_{N})$, the mean and covariance functions are given by
@@ 2.1 Basis decomposition
$\mu(\mathbf{t})=\Psi(\mathbf{t})^{\top}\mathbf{C}^{\top}\Pi\quad\text{and}\quad C(\mathbf{s},\mathbf{t})=\Psi(\mathbf{s})^{\top}\mathbf{C}^{\top}\left(\text{diag}\{\pi_{1},\dots,\pi_{N}\}-\Pi\Pi^{\top}\right)\mathbf{C}\Psi(\mathbf{t}).$ (12)
Finally, we denote by $\mathbf{W}$ the matrix of inner products of the functions in the basis $\Psi$. The matrix $\mathbf{W}$ is a block-diagonal matrix such that $\mathbf{W}=\text{diag}\{\mathbf{W}^{(1)},\dots,\mathbf{W}^{(P)}\}$ where each entry is given by
$\mathbf{W}_{k,l}^{(p)}=\left\langle\psi_{k}^{(p)},\psi_{l}^{(p)}\right\rangle,\quad 1\leq k,l\leq K_{p},\quad 1\leq p\leq P.$ (13)
We remark that, if the basis $\Psi$ is an orthonormal basis, the matrix $\mathbf{W}$ is equal to the identity matrix of size $\sum_{p=1}^{P}K_{p}$. Using the expansion of the data into the basis of functions $\Psi$, the inner-product matrix $\mathbf{M}$ is written
$\mathbf{M}=\text{diag}\{\sqrt{\pi_{1}},\dots,\sqrt{\pi_{N}}\}\left(\mathrm{I}_{\!N}-\mathbf{1}_{\!N}\Pi^{\top}\right)\mathbf{C}\mathbf{W}\mathbf{C}^{\top}\left(\mathrm{I}_{\!N}-\Pi\mathbf{1}_{\!N}^{\top}\right)\text{diag}\{\sqrt{\pi_{1}},\dots,\sqrt{\pi_{N}}\}$ (14)
where $\mathrm{I}_{\!N}$ is the identity matrix of size $N$ and $\mathbf{1}_{\!N}$ is a vector of $1$ of length $N$.
@@ 3.1 Duality diagram
The distinction between the space of rows of a matrix as a sample from a population and the space of columns as the fixed variables on which the observations were measured has been explained in Holmes, (2008) and De la Cruz and Holmes, (2011) for multivariate data analysis. We propose to define a duality diagram in the context of multivariate functional data. Consider the data matrix defined by the set $\mathcal{X}$. We define an operator $L_{X}:\mathcal{H}\rightarrow\mathbb{R}^{N}$ by
@@
Then $L^{\star}_{X}$ is the adjoint operator of the linear operator $L_{X}$ (see Appendix A for a proof). As an adjoint operator, $L^{\star}_{X}$ is a linear operator. The operator $\Gamma$ and the matrix $\mathbf{M}$ define geometries in $\mathcal{H}$ and $\mathbb{R}^{N}$, respectively, through
@@ 3.2 Cloud of individuals
The utilization of the distance measure $d_{\Gamma}$, which accounts for the variability among all the features within the functional data, corresponds to a Mahalanobis-type distance framework for multivariate functional data (see Berrendero et al., (2020) and Martino et al., (2019)). Given the set $\mathcal{X}$, the total inertia of $\mathcal{C}_{\!P}$, with respect to $\mathrm{G}_{\!\mu}$ and the distance $d_{\Gamma}$, is given by
@@ 3.3 Cloud of features
Given an element $f\in\mathcal{H}$, let $L_{X}(f)=\{\sqrt{\pi_{n}}\langle\!\langle X_{n}-\mu,f\rangle\!\rangle,\,n=1,\dots,N\}$ be the set of projections of $f$ onto the centered observations. We identify this set as the point $\mathsf{M}_{f}$ in the space $\mathbb{R}^{N}$. The space $\mathbb{R}^{N}$ is referred to as the features’ space. The cloud of points that represents the set of observations in $\mathbb{R}^{N}$ is denoted by $\mathcal{C}_{\!N}$. Let $\mathsf{G}_{\!\mu}$ be the centre of gravity of the cloud $\mathcal{C}_{\!N}$. In the space $\mathbb{R}^{N}$, its coordinates are given by $L_{X}(\mu)=\{\sqrt{\mu}\langle\!\langle X_{n}-\mu,\mu\rangle\!\rangle,\,n=1,\dots,N\}$. If the data are centered, the origin $\mathsf{O}_{\!\mathbb{R}}$ of the axes in $\mathbb{R}^{N}$ coincides with $\mathsf{G}_{\!\mu}$.

### v2

# supplement

@@ 1 Introduction
The key motivation of this paper is to investigate the duality between rows and columns of a data matrix to estimate the eigencomponents of a multivariate functional dataset. The duality between rows and columns of a data matrix is a fundamental concept in classical multivariate statistics (Escofier,, 1979; Saporta,, 1990). A data matrix typically represents a set of observations of multiple features, each row corresponds to an individual observation and each column corresponds to an individual feature. The duality between rows and columns refers to the fact that many statistical methodologies can be conducted either on the rows or the columns of the data matrix, and the results will be related to each other. For example, the principal components obtained from a PCA run on the rows of the data matrix are the same as the ones obtained from a PCA run on the columns of the matrix. The choice of method to use, based on criteria such as computational time or data storage needs, is thus left to the user. This concept has been widely studied in multivariate statistics (see, e.g., Pagès, (2014); Härdle and Simar, (2019)). In the context of functional data, this principle has received limited attention despite being mentioned in the seminal paper of FDA (Ramsay,, 1982). Ramsay and Silverman, (2005) briefly commented on it in a concluding remark of Chapter 8, while Kneip and Utikal, (2001) and Benko et al., (2009) utilized it to compute principal components for dense univariate functional data. Chen et al., (2017) also employ it to gain computational advantage when univariate functional data are sampled on a very dense grid. To the best of our knowledge, however, there is no available literature on its application to multivariate functional data that are observed on different dimensional domains. Our aim is therefore to investigate this duality for multivariate functional data observed on different dimensional domains and provide guidelines on which method to use in different cases.
The remainder of the paper is organized as follows. In Section 2, we define multivariate functional data with components that are observed on possibly different domains. In Section 3, we develop the duality between the observations’ space and the functional components’ space. The relationship between the eigencomponents of the covariance operator of the multivariate functional datasets and the eigencomponents of the inner-product matrix between the observations is derived in Section 4. Extensive simulations are given in Section 5. We also provide guidelines on which method to use with respect to different data characteristics. We present an application related to sports science data from the National Basketball Association (NBA) in Section 6. The paper concludes with a discussion and an outlook in Section 7.
@@ 2 Model
Let us consider a set of $N$ independent multivariate curves $\mathcal{X}=\{X_{n}\}_{1\leq n\leq N}$ generated as a random sample of the $P$-dimensional stochastic process $X$ with continuous trajectories. The data can be viewed as a table with $N$ rows and $P$ columns where each entry is a curve, potentially on a multidimensional domain (see Figure 1). Each row of this matrix represents an observation; while each column represents a functional feature. At the intersection of row $n$ and column $p$, we thus have $X_{n}^{(p)}$ which is the curve that concerns the (functional) feature $p$ for the individual $n$. For $n\in\{1,\dots,N\}$, each observation $n$ is attributed the weight $\pi_{n}$ such that $\sum_{n}\pi_{n}=1$, e.g., $\pi_{n}=1/N$. For the set $\mathcal{X}$, the inner-product matrix, also called the Gram matrix, $\mathbf{M}$ is defined as a matrix of size $N\times N$ with entries
This matrix is symmetric, positive definite, and interpretable as a proximity matrix, each entry being the similarity between the weighted observations.
@@
This setting can be generalized to incorporate a general weighting scheme, as Chiou et al., (2014) or Happ and Greven, (2018). We consider the function $\langle\!\langle\cdot,\cdot\rangle\!\rangle_{w}~{}:~{}\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$,
$\langle\!\langle f,g\rangle\!\rangle_{w}\coloneqq\sum_{p=1}^{P}w_{p}\left\langle f^{(p)},g^{(p)}\right\rangle=\sum_{p=1}^{P}w_{p}\int_{\mathcal{T}_{p}}f^{(p)}(t_{p})g^{(p)}(t_{p}){\rm d}t_{p},\quad f,g\in\mathcal{H}.$ (6)
Let $\Gamma_{w}:\mathcal{H}\rightarrow\mathcal{H}$ denote the weighted covariance operator of $X$, defined as an integral operator with kernel $C$. That is, for $f\in\mathcal{H}$ and $\mathbf{t}\in\mathcal{T}$, the $p$th feature of $\Gamma_{w}f(\mathbf{t})$ is given by
$(\Gamma_{w}f)^{(p)}(t_{p})\coloneqq\langle\!\langle C_{p,\cdot}(t_{p},\cdot),f(\cdot)\rangle\!\rangle_{w}=\langle\!\langle C_{\cdot,p}(\cdot,t_{p}),f(\cdot)\rangle\!\rangle_{w},\quad t_{p}\in\mathcal{T}_{p}.$ (7)
@@
Similarly, we define the weigthed inner-product matrix $\mathbf{M}_{w}$, the matrix of size $N\times N$ with entries
$\mathbf{M}_{w,nn^{\prime}}=\sqrt{\pi_{n}\pi_{n^{\prime}}}\langle\!\langle X_{n}-\mu,X_{n^{\prime}}-\mu\rangle\!\rangle_{w},\quad n,n^{\prime}=1,\dots,N.$ (8)
Remark 1.
The observation weights $\pi_{n}$ and the feature weights $w_{p}$ are not the same and should not be confused. The observation weights are used to give more (or less) importance to specific observations, while the feature weights control for different levels of variation for the features. A discussion on the feature weights is provided in Section 3.4.
2.1 Inference
If we observe the set $\mathcal{X}$, the ideal estimators of the mean and covariance function are
$\widetilde{\mu}(\mathbf{t})=\sum_{n=1}^{N}\pi_{n}X_{n}(\mathbf{t}),\quad\text{and}\quad\widetilde{C}(\mathbf{s},\mathbf{t})=\sum_{n=1}^{N}\pi_{n}\{X_{n}(\mathbf{s})-\widetilde{\mu}(\mathbf{s})\}\{X_{n}(\mathbf{t})-\widetilde{\mu}(\mathbf{t})\}^{\top},\quad\mathbf{s},\mathbf{t}\in\mathcal{T}.$ (9)
Similarly, one could estimate the inner-product $\mathbf{M}$ by replacing the mean by their estimators. The resulting matrix $\widetilde{\mathbf{M}}$ has entries
$\widetilde{\mathbf{M}}_{nn^{\prime}}=\sqrt{\pi_{n}\pi_{n^{\prime}}}\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\{X_{n}^{(p)}(t_{p})-\widetilde{\mu}^{(p)}(t_{p})\}\{X_{n^{\prime}}^{(p)}(t_{p})-\widetilde{\mu}^{(p)}(t_{p})\}{\rm d}t_{p},\quad n,n^{\prime}=1,\dots,N.$ (10)
In many applications, the elements of the set $\mathcal{X}$ are observed with error and on a finite grid of points in $\mathcal{T}$. For each $1\leq n\leq N$, and given a vector of positive integers $\mathsf{M}_{n}=(\mathsf{M}_{n}^{(1)},\dots,\mathsf{M}_{n}^{(P)})$, let $\mathrm{T}_{n,\mathsf{m}}=(\mathrm{T}_{n,m_{1}}^{(1)},\dots,\mathrm{T}_{n,m_{P}}^{(P)}),1\leq m_{p}\leq\mathsf{M}_{n}^{(p)},1\leq p\leq P$, be the random observation times for the curve $X_{n}$. These times are obtained as independent realizations of a random variable $\mathbf{T}$ taking values in $\mathcal{T}$. The vectors $\mathsf{M}_{1},\dots,\mathsf{M}_{N}$ represent an independent sample of an integer-valued random vector $\mathsf{M}$ with known expectation. We assume that the realizations of $X$, $\mathsf{M}$ and $\mathbf{T}$ are mutually independent. The observations associated with an observation $X_{n}$ consist of the pairs $(Y_{n,\mathsf{m}},\mathrm{T}_{n,\mathsf{m}})\in\mathbb{R}^{P}\times\mathcal{T}$, where $\mathsf{m}=(m_{1},\dots,m_{P}),1\leq m_{p}\leq\mathsf{M}_{n}^{(p)},1\leq p\leq P$ and $Y_{n,\mathsf{m}}$ is defined as
$Y_{n,\mathsf{m}}=X_{n}(\mathrm{T}_{n,\mathsf{m}})+\varepsilon_{n,\mathsf{m}},\quad 1\leq n\leq N,$ (11)
with the $\varepsilon_{n,\mathsf{m}}$ being independent realizations of a centered error random vector $\varepsilon\in\mathbb{R}^{P}$ with finite variance $\sigma^{2}\mathbf{1}_{P}\in\mathbb{R}^{P\times P}$ where $\sigma^{2}=(\sigma^{2}_{1},\dots,\sigma^{2}_{P})\in\mathbb{R}^{P}$.
Let $\widehat{X}_{n}$ be a suitable estimator of $X_{n}$ applied to the pairs $(Y_{n,\mathsf{m}},\mathrm{T}_{n,\mathsf{m}})$, such as P-splines (e.g. Eilers et al.,, 2015) or local polynomials (e.g. Fan and Gijbels,, 1996). We define the estimator of the mean function as
$\widehat{\mu}(\mathbf{t})=\sum_{n=1}^{N}\pi_{n}\widehat{X}_{n}(\mathbf{t}),\quad\mathbf{t}\in\mathcal{T}.$ (12)
Concerning the estimation of the covariance of the $p$th feature, we distinguish the diagonal from the non-diagonal points. The estimation of the covariance function of the non-diagonal points is defined as
$\widehat{C}_{p,p}(s_{p},t_{p})=\sum_{n=1}^{N}\pi_{n}\left(\{\widehat{X}_{n}^{(p)}(s_{p})-\widehat{\mu}^{(p)}(s_{p})\}\{\widehat{X}_{n}^{(p)}(t_{p})-\widehat{\mu}^{(p)}(t_{p})\}\right),\quad s_{p}\neq t_{p},\quad s_{p},t_{p}\in\mathcal{T}_{p}.$ (13)
The variance function $C_{p,p}(s_{p},t_{p})$ induces a singularity when estimating the covariance function $C_{p,p}(\cdot,\cdot)$ (see Yao et al.,, 2005; Zhang and Wang,, 2016). We define the estimator of the diagonal of the covariance as
$\widehat{C}_{p,p}(s_{p},s_{p})=\sum_{n=1}^{N}\pi_{n}\{\widehat{X}_{n}^{(p)}(s_{p})-\widehat{\mu}^{(p)}(s_{p})\}^{2}-\sigma_{p}^{2},\quad s_{p}\in\mathcal{T}_{p}.$ (14)
This singularity does not appear in the estimation of the cross-covariance function. We define the estimator of the cross-covariance function between the $p$th and $q$th features by
$\widehat{C}_{p,q}(s_{p},t_{q})=\sum_{n=1}^{N}\pi_{n}\left(\{\widehat{X}_{n}^{(p)}(s_{p})-\widehat{\mu}^{(p)}(s_{p})\}\{\widehat{X}_{n}^{(q)}(t_{q})-\widehat{\mu}^{(q)}(t_{q})\}\right),\quad s_{p}\in\mathcal{T}_{p},t_{q}\in\mathcal{T}_{q}.$ (15)
The estimator of the inner-product matrix can be defined by replacing the curves by their estimators. Following Benko et al., (2009) and Grith et al., (2018), we define the estimator of the inner-product matrix as
$\widehat{\mathbf{M}}_{nn^{\prime}}=\sqrt{\pi_{n}\pi_{n^{\prime}}}\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\{\widehat{X}_{n}^{(p)}(t_{p})-\widehat{\mu}^{(p)}(t_{p})\}\{\widehat{X}^{(p)}_{n^{\prime}}(t_{p})-\widehat{\mu}^{(p)}(t_{p})\}{\rm d}t_{p},\quad n,n^{\prime}=1,\dots,N.$ (16)
As for the covariance estimation, the model implies a bias on the diagonal term of the inner-product matrix. A correction for the diagonal of the Gram matrix is given by
@@ 2.1 Inference
$\widehat{\mathbf{M}}_{nn}=\pi_{n}\sum_{p=1}^{P}\left\{\int_{\mathcal{T}_{p}}\{\widehat{X}_{n}^{(p)}(t_{p})-\widehat{\mu}^{(p)}(t_{p})\}^{2}{\rm d}t_{p}-\sigma_{p}^{2}\right\},\quad n=1,\dots,N.$ (17)
For the estimation of the variances $\sigma_{p}^{2}$, we let the reader refers to Hall and Marron, (1990) and Hall et al., (1990).
@@ 3.1 Duality diagram
The distinction between the space of rows of a matrix as a sample from a population and the space of columns as the fixed variables on which the observations were measured has been explained in Holmes, (2008) and De la Cruz and Holmes, (2011) for multivariate data. We propose to define a duality diagram in the context of multivariate functional data. Consider the data matrix defined by the set $\mathcal{X}$. We define an operator $L_{X}:\mathcal{H}\rightarrow\mathbb{R}^{N}$ by
@@
Lemma 1.
$L^{\star}_{X}$ is the adjoint operator of the linear operator $L_{X}$.
The proof of Lemma 1 is given in Appendix A. As an adjoint operator, $L^{\star}_{X}$ is a linear operator. The operator $\Gamma$ and the matrix $\mathbf{M}$ define geometries in $\mathcal{H}$ and $\mathbb{R}^{N}$, respectively, through
@@ 3.2 Cloud of individuals
The utilization of the distance measure $d_{\Gamma}$, which accounts for the variability among all the features within the functional data, corresponds to a Mahalanobis-type distance framework for multivariate functional data (see Berrendero et al.,, 2020; Martino et al.,, 2019). Given the set $\mathcal{X}$, the total inertia of $\mathcal{C}_{\!P}$, with respect to $\mathrm{G}_{\!\mu}$ and the distance $d_{\Gamma}$, is given by
@@ 3.3 Cloud of features
Given an element $f\in\mathcal{H}$, let $L_{X}(f)=\{\sqrt{\pi_{n}}\langle\!\langle X_{n}-\mu,f\rangle\!\rangle,\,n=1,\dots,N\}$ be the set of projections of $f$ onto the centered observations. We identify this set as the point $\mathsf{M}_{f}$ in the space $\mathbb{R}^{N}$. The space $\mathbb{R}^{N}$ is referred to as the features’ space. The cloud of points that represents the set of observations in $\mathbb{R}^{N}$ is denoted by $\mathcal{C}_{\!N}$. Let $\mathsf{G}_{\!\mu}$ be the centre of gravity of the cloud $\mathcal{C}_{\!N}$. In the space $\mathbb{R}^{N}$, its coordinates are given by $L_{X}(\mu)=\{\sqrt{\pi_{n}}\langle\!\langle X_{n}-\mu,\mu\rangle\!\rangle,\,n=1,\dots,N\}$. If the data are centered, the origin $\mathsf{O}_{\!\mathbb{R}}$ of the axes in $\mathbb{R}^{N}$ coincides with $\mathsf{G}_{\!\mu}$.